# HRET: A Self-Evolving LLM Evaluation Toolkit for Korean

### v2

@@ Abstract
Recent advancements in Korean large language models (LLMs) have spurred numerous benchmarks and evaluation methodologies, yet the lack of a standardized evaluation framework has led to inconsistent results and limited comparability. To address this, we introduce HRET (Haerae Evaluation Toolkit), an open-source, self-evolving evaluation framework tailored specifically for Korean LLMs. HRET unifies diverse evaluation methods, including logit-based scoring, exact-match, language-inconsistency penalization, and LLM-as-a-Judge assessments. Its modular, registry-based architecture integrates major benchmarks (HAE-RAE Bench, KMMLU, KUDGE, HRM8K) and multiple inference backends (LiteLLM, HuggingFace, OpenAI-compatible endpoints). With automated pipelines for continuous evolution, HRET provides a robust foundation for reproducible, fair, and transparent Korean NLP research.
HRET: A Self-Evolving LLM Evaluation Toolkit for Korean
Hanwool Lee1,2  Sooyong Kim1  Dasol Choi1,3  Sangwon Baek1,4 Seunghyeok Hong1,5  Ilgyun Jeong1,6  Naeun Lee1  Inseon Hwang1  Guijin Son1,7 1HAERAE 2Shinhan Securities  3Yonsei University  4Catius  5Hankuk University of Foreign Studies 6Korea University  7OneLineAI gksdnf424@gmail.com
@@ 1 Introduction
Mature evaluation practices are essential for accurately tracking advancements in the capabilities of large language models (LLMs). Biased or skewed assessments can mislead the research community and hinder shared goals of technological progress. In this line, the English NLP community has long pursued credible, standardized benchmarks—starting with the GLUE and SuperGLUE leaderboards (Wang et al., 2018, 2019), followed by LM-Eval Harness (Gao et al., 2024), HELM (Liang et al., 2023), and Chatbot Arena—all aimed at providing fair, comparable evaluations under consistent conditions. Now, as Korean NLP research increases in volume, they face similar challenges: Technical reports for foundational models often yield inconsistent metrics, complicating direct comparisons. This underscores the need for a unified evaluation framework tailored to Koreans.
Korean benchmarks vary considerably in their evaluation setups—HAE-RAE Bench utilizes a log-probability-based evaluation, while KMMLU employs exact-match criteria. Such methodological discrepancies often compel researchers and practitioners to customize evaluations, complicating direct comparisons and compromising interpretability. For example, when evaluating identical models like Gemma2-27B on the same benchmarks, different research groups report significantly different results (as shown in Table 1), highlighting the urgent need for standardization.
Additionally, existing evaluations frequently overlook key linguistic nuances unique to Korean, such as intricate morphology, context-dependent word segmentation, and extensive honorific systems. These language-specific challenges require specialized consideration in evaluation methodologies to ensure accurate performance assessment of Korean LLMs.
To address these challenges, we introduce HRET (Haerae Evaluation Toolkit), a standardized, open-source toolkit specifically tailored for evaluating Korean LLMs. HRET makes several key contributions:
  • Unified and reproducible evaluation: HRET standardizes dataset handling, evaluation methods, and benchmarking procedures, ensuring transparency and consistent outcomes.
  • Self-Evolving Mechanism: Its modular registry-based design allows HRET to continuously integrate new benchmarks, evaluation methodologies, and inference backends with minimal effort.
  • Multi-method evaluation: Supports multiple complementary evaluation methods—including logit-based scoring, exact-match evaluation, and LLM-as-a-judge—to comprehensively assess model performance.
  • Ease of Use: Intuitive API and streamlined command-line interfaces simplify evaluation workflows, making it accessible for both researchers and industry practitioners.
By unifying prominent benchmarks and diverse inference backends—including HuggingFace, vLLM, and OpenAI-compatible APIs—HRET significantly improves reproducibility, comparability, and accessibility in Korean NLP research. The toolkit and comprehensive documentation will be publicly released under the Apache License 2.0 upon acceptance to encourage widespread adoption and community contributions.
@@
Table 1: Performance discrepancies reported for the Gemma-2-27B model between EXAONE 3.5 and Kanana technical reports, highlighting reproducibility challenges in Korean NLP evaluations.
@@ 2.2 Reproducibility Challenges
Current Korean NLP evaluations exhibit considerable reproducibility challenges. A notable example is the inconsistent results reported for widely used models such as Gemma2-27B Team et al. (2024) and Qwen2.5-32B Yang et al. (2024) across various technical reports from major Korean organizations. As illustrated in Table 1, evaluations of identical benchmarks using these same models yield significantly different outcomes depending on the evaluating institution. For instance, the EXAONE 3.5 report Research et al. (2024) and Kanana report Bak et al. (2025) report performance discrepancies ranging from around 1–10 percentage points across multiple tasks, including GSM8K, KMMLU, HumanEval, and MBPP.
These discrepancies stem from variations in inference settings, prompt engineering, and evaluation implementations. The situation is further complicated by Korean’s linguistic features—complex morphology, context-dependent segmentation, and honorific systems—which are handled inconsistently across evaluations. Existing benchmarks lack standardized guidelines addressing these Korean-specific challenges, resulting in fragmented evaluation practices that impede meaningful progress in Korean NLP research.
@@
Developer EleutherAI OpenAI UpstageAI Confident AI HAERAE
@@
Evaluation Metrics Traditional metrics Task-specific scripts Integrated metrics 14+ diverse metrics Traditional + LLM-as-judge
@@
Automation Script-based API-based No-code automation CI/CD integration Pipeline Runner
Unique Strength High reproducibility OpenAI optimization Multi-tool integration Testing-focused Registry-based extensibility
@@ 2.3 Comparison with Existing Systems
HRET differentiates itself through its specific focus on Korean language evaluation challenges, its registry-based architecture enabling easy extension, and its integrated pipeline approach that combines dataset handling, model inference, scaling, and evaluation into a unified framework.
@@ 3.1 Novelty and Design Principles
  • Self-Evolving Mechanism: HRET continuously adapts by automatically discovering and integrating new evaluation methods, datasets, and inference backends at runtime through (1) Automated Component Discovery for runtime integration without core code modification, (2) Dynamic Pipeline Construction that adapts to new scenarios without predefined workflows, and (3) Continuous Integration Workflows that validate components against established benchmarks.
  • Registry-Based Modularity: Core components—including dataset loaders, evaluation methods, scaling strategies, and inference backends—are individually modularized and managed through registries, allowing researchers to effortlessly extend the toolkit’s capabilities with minimal code changes.
@@
  • Korean Focus: The framework specifically targets Korean language benchmarks and evaluation needs, with built-in support for major Korean evaluation datasets and specialized evaluation methods like language consistency checking.
Category Benchmark Description
General HAE-RAE Son et al. (2023) Evaluation of Korean language understanding, focusing on vocabulary knowledge, reading comprehension, and general knowledge across multiple domains.
KMMLU Son et al. (2024b) Assessment of multi-domain knowledge across 45 categories including STEM, humanities, applied sciences, and other fields in Korean language context.
CLIcK Kim et al. (2024a) Measurement of cultural and linguistic proficiency through 1,995 QA pairs from Korean exams and textbooks, organized into eleven distinct categories.
Dialogue KUDGE Son et al. (2024a) Evaluation of dialogue generation capabilities, with emphasis on coherence and contextual appropriateness in Korean conversational settings.
Specialized K2-Eval Son et al. (2024a) Assessment of domain-specific knowledge in professional, technical, and academic fields requiring specialized Korean terminology.
HRM8K Ko et al. (2025) Evaluation of mathematical reasoning using 8,011 problems, combining 1,428 native Korean math exam questions and 6,583 translated problems from other benchmarks.
Table 3: Descriptions of benchmarks supported in HRET.
@@ 3.2 System Architecture and Components
HRET provides a unified interface for evaluating major Korean NLP benchmarks (detailed in Table 3), following a standardized five-stage workflow:
@@
    Dataset Loading: The Dataset Registry unifies Korean NLP datasets with consistent formatting, supporting standard benchmarks and offering a flexible generic_file loader for custom datasets without requiring dedicated classes.
@@
    Model Loading: The Backend Registry connects to various backends (HuggingFace, OpenAI-compatible, LiteLLM), providing a unified interface that abstracts provider-specific details while supporting generation, judge, and reward functionalities.
@@
    Inference with Scaling: Scaling Techniques enhance evaluation quality through (1) Beam Search  (Chen et al., 2024) for exploring different decoding paths, (2) Best-of-N Sampling  (Stiennon et al., 2022) for selecting optimal outputs, and (3) Majority Voting  (Wang et al., 2023) for enhancing stability through consensus.
@@
    Evaluation: Generated predictions are assessed using multiple approaches: (1) String-matching for reference comparison, (2) Logit-based Evaluation for confidence analysis, and (3) LLM-as-a-Judge for open-ended assessment. Additional Evaluation Enhancements include Language penalizer for detecting non-Korean responses and Answer extraction for identifying final answers from reasoning chains.
@@
3.3 Advanced Features and Usage
HRET supports multiple inference backends through a consistent interface:
  • HuggingFace Transformers: Provides direct access to thousands of models, including Korean-specific models like KoGPT and HyperCLOVA.
  • vLLM: Optimized for high-throughput evaluation scenarios with large Korean benchmarks.
  • OpenAI-compatible endpoints: Ensures compatibility with API-based services from major Korean AI providers.
  • LiteLLM: Offers unified access to multiple model providers, simplifying cross-vendor comparisons.
Beyond the standard Evaluator API, HRET provides a flexible PipelineRunner for advanced customization of evaluation workflows. This interface enables:
  • Multi-stage evaluations where outputs from one model serve as inputs to another
  • Comprehensive model comparisons across different benchmarks and metrics
  • Custom workflows with specialized processing steps for Korean-specific evaluations
  • Ensemble evaluations combining complementary models or methodologies
@@ 3.4 Usage Examples
To illustrate the ease of use and practical value of HRET, we provide a brief usage example evaluating a Korean LLM on the HAE-RAE Bench benchmark using our command-line interface (CLI):
@@
This command: (1) Loads the specified benchmark dataset (‘HAE-RAE Bench‘, subset ‘csat_geo‘), (2) Instantiates the chosen HuggingFace model (‘Gemma-3-1b-it‘), (3) Evaluates the model outputs using the exact-match criteria, and (4) Outputs evaluation results in JSON format.
Additional detailed usage scenarios and advanced functionalities (Python API, multi-method evaluations, scaling methods) are provided in Appendix A.
4 System Validation and Roadmap
4.1 QA Testing and Benchmarking
HRET underwent rigorous validation to ensure reliability and usability:
QA Testing
We conducted extensive functional testing through a rigorous QA process, ensuring all components performed reliably under diverse configurations and use cases.
Real-World Deployment
The framework was deployed as the backend system for the HAE-RAE Reasoning Challenge leaderboard, demonstrating its robustness in a production environment with various model submissions and evaluation criteria.
Performance Benchmarking
We conducted extensive performance testing to ensure the system maintains efficiency with large-scale Korean benchmark datasets, validating that the asynchronous and multi-threading capabilities deliver the expected throughput improvements.
4.2 Current Status and Roadmap
HRET is under active development, currently supporting all benchmarks mentioned and the evaluation methodologies described. The ongoing roadmap includes strengthening the self-evolving mechanism through enhanced implementation of the hret-agent, which leverages the plugin-based registry architecture for improved scalability. Additionally, HRET aims to deepen integration and interoperability with other open-source LLM libraries, promoting broader adoption and community collaboration.
5 Experiments
We conduct two experiments to highlight how HRET effectively addresses critical challenges specific to Korean LLM evaluations.
5.1 Language Consistency Evaluation with Language Penalizer
Multilingual models often respond in non-Korean languages, particularly English, skewing evaluations on Korean tasks. HRET’s language penalizer detects and penalizes such responses, ensuring results accurately reflect Korean language capabilities. Figure 4 shows significant performance adjustments for several models on the K2-Eval benchmark when language penalization is enabled.
Figure 2: Performance comparison with and without language penalizer on K2-Eval. Qwen models show reductions when enforced, while EXAONE and Llama models maintain consistent scores.
5.2 Multi-Method Evaluation Approaches
HRET supports multiple evaluation methods tailored for different task types. Table 4 compares the performance of EXAONE-3.5-2.4B across evaluation methods on HRM8K and CLIcK benchmarks.
Benchmark String-Match Math-Verify
HRM8K 2.21 14.11
CLIcK 27.31 28.67
@@ 5.2 Multi-Method Evaluation Approaches
Table 4: Comparison of evaluation methods for EXAONE-3.5-2.4B across benchmarks.
Key insights include:
  • Performance varies significantly with evaluation method, particularly for mathematical reasoning tasks (HRM8K shows a 6.4x improvement with Math-Verify).
  • For language understanding datasets like CLIcK, evaluation methods yield more consistent results (difference of only 1.36%).
  • Task-appropriate evaluation metrics provide more accurate assessments of model capabilities.
  • HRET enables researchers to select the most suitable evaluation method for different Korean language tasks.
These experiments demonstrate how HRET’s diverse evaluation methodologies reveal important performance nuances that might be missed with a single evaluation approach.

### v3

# Redefining Evaluation Standards: A Unified Framework for Evaluating the Korean Capabilities of Language Models

@@ Abstract
Recent advancements in Korean large language models (LLMs) have driven numerous benchmarks and evaluation methods, yet inconsistent protocols cause up to 10 pp performance gaps across institutions. Overcoming these reproducibility gaps does not mean enforcing a one-size-fits-all evaluation. Rather, effective benchmarking requires diverse experimental approaches and a framework robust enough to support them. To this end, we introduce HRET (Haerae Evaluation Toolkit), an open-source, registry-based framework that unifies Korean LLM assessment. HRET integrates major Korean benchmarks, multiple inference backends, and multi-method evaluation, with language consistency enforcement to ensure genuine Korean outputs. Its modular registry design also enables rapid incorporation of new datasets, methods, and backends, ensuring the toolkit adapts to evolving research needs. Beyond standard accuracy metrics, HRET incorporates Korean-focused output analyses—morphology-aware Type–Token Ratio (TTR) for evaluating lexical diversity and systematic keyword-omission detection for identifying missing concepts—to provide diagnostic insights into language-specific behaviors. These targeted analyses help researchers pinpoint morphological and semantic shortcomings in model outputs, guiding focused improvements in Korean LLM development.
Redefining Evaluation Standards: A Unified Framework for Evaluating the Korean Capabilities of Language Models
Hanwool Lee1,2thanks: Equal contribution., Dasol Choi1,311footnotemark: 1, Sooyong Kim1, Ilgyun Jung1,4, Sangwon Baek1,6, Guijin Son1,7, Inseon Hwang1, Naeun Lee1, Seunghyeok Hong1,5thanks: Corresponding author. 1MODULABS, HAERAE LAB 2Shinhan Securities  3Yonsei University  4Catius  5Hankuk University of Foreign Studies 6Korea University  7OneLineAI gksdnf424@gmail.com
@@ 1 Introduction
Recent advances in large language models (LLMs) demand rigorous, reproducible evaluation, yet real-world benchmarking often diverges from official protocols Liang et al. (2022); Gao et al. (2024). In Korean NLP, identical models have exhibited performance gaps up to 10 percentage points across institutions Research et al. (2024); Bak et al. (2025), driven by variations in prompt templates, inference settings, and evaluation criteria. We argue that true reproducibility arises not from a rigid framework but from a toolkit that flexibly supports multiple evaluation paradigms—few-shot versus log-likelihood scoring, exact-match versus generative judgment, and test-time scaling techniques—so that assessments can adapt to both model characteristics and deployment contexts.
Consider KMMLU Son et al. (2024b), whose official log-likelihood/direct-generation implementation in lm-eval-harness is rarely used unchanged: practitioners often augment it with n-shot in-context examples Brown et al. (2020), inject chain-of-thought prompts, or apply self-consistency Wang et al. (2023) to better capture model reasoning. Such customization highlights the need for an extensible evaluation framework that integrates diverse Korean benchmarks and inference backends while exposing configurable strategies for prompt design, scoring, and aggregation.
To meet this need, we present HRET (Haerae Evaluation Toolkit), an open-source, registry-based framework for unified Korean LLM assessment. HRET standardizes data ingestion and result reporting, supports string matching, partial matching, log-likelihood, mathematical verification, and LLM-as-a-judge methods Zheng et al. (2023), and applies a language-penalty mechanism to discourage non-Korean outputs. Importantly, it embeds Korean-specific diagnostics—morphology-aware Type–Token Ratio for lexical diversity and systematic keyword-omission detection—to surface linguistic phenomena unique to Korean. Its modular registry design enables rapid onboarding of new datasets, methods, and backends, future-proofing evaluation as the field evolves.
We validate HRET’s robustness and efficiency across diverse hardware configurations, demonstrating substantial reductions in evaluation variability and improved throughput. HRET thus offers a practical solution for reproducible, adaptable benchmarking in Korean LLM research.
@@
Table 1: Performance discrepancies for identical models Qwen et al. (2025); Team et al. (2024) between EXAONE 3.5 and Kanana technical reports, highlighting reproducibility challenges in Korean NLP evaluations.
@@ 2.2 Reproducibility Challenges
Evaluations of Korean language models often show inconsistent results across institutions, with the same models reporting performance gaps of up to 10 percentage points. For instance, Gemma2-27B and Qwen2.5-32B Qwen et al. (2025) achieve scores that differ by 1–10 pp on GSM8K, KMMLU, HumanEval and MBPP when comparing the EXAONE 3.5 report Research et al. (2024) and the Kanana report Bak et al. (2025) (Table 1). This variability underscores the absence of consistent evaluation practices tailored to Korean LLMs, complicating direct comparisons and hindering reproducibility.
These gaps arise from varying inference settings, prompt designs, and evaluation implementations. On top of that, Korean’s own characteristics—its rich morphology, context-dependent tokenization, and honorific system—make consistent handling difficult without standardized, Korean-focused guidelines. Together, these factors exacerbate result variability and impede reliable benchmarking across research groups.
@@
Developer EleutherAI OpenAI Upstage Confident AI HAERAE
@@
Evaluation Metrics Traditional metrics Task-specific scripts Integrated metrics 5+ diverse metrics 5+ diverse metrics
@@
Automation Script-based API-based No-code automation CI/CD integration API-based, CI/CD integration
Unique Strength High reproducibility OpenAI optimization Multi-tool integration Testing-focused Korean-specific + Registry extensibility
@@ 2.3 Comparison with Existing Systems
HRET differentiates itself through three key aspects: (1) Korean-specific capabilities including language consistency enforcement and concise morphological analysis–based output diagnostics, (2) registry-based extensibility that enables rapid component integration compared to traditional monolithic architectures, and (3) unified evaluation pipeline that unifies fragmented Korean LLM assessment practices. Unlike existing frameworks that focus on general-purpose evaluation, HRET addresses specific reproducibility challenges observed in Korean NLP research (Table 1).
@@ 3.1 Design Principles
  • Registry-Based Modularity: Core components—including dataset loaders, evaluation methods, scaling strategies, and inference backends—are individually modularized and managed through registries, allowing researchers to effortlessly extend the toolkit’s capabilities with minimal code changes compared to traditional monolithic architectures.
  • Unified Standardization: HRET systematically standardizes fragmented Korean LLM evaluation practices by providing consistent interfaces, reproducible workflows, and transparent benchmarking procedures across different research environments.
  • Korean Language Consistency: Penalizes any response not in Korean by applying a score penalty, ensuring evaluation focuses strictly on Korean outputs.
@@ 3.2 System Architecture and Components
HRET provides a unified interface for evaluating major Korean NLP benchmarks including HAE-RAE Bench, KMMLU, KUDGE, HRM8K, CLIcK, and others (see Appendix A for complete list), following a standardized five-stage workflow:
@@
    Dataset Loading: The Dataset Registry unifies Korean NLP datasets with consistent formatting, supporting standard benchmarks and offering a flexible generic_file loader for custom datasets without dedicated classes.
@@
    Model Loading: The Backend Registry connects to various backends (HuggingFace, vLLM, OpenAI-compatible, etc), providing a unified interface that abstracts provider-specific details while supporting generation, judge, and reward functionalities.
@@
    Inference with Scaling: Scaling Techniques enhance evaluation quality through (1) Beam Search for exploring different decoding paths, (2) Best-of-N Sampling for selecting optimal outputs, and (3) Majority Voting for enhancing stability through consensus.
@@
    Evaluation: Generated predictions are assessed using multiple approaches: (1) String-matching, (2) Logit-based Evaluation (3) Mathematical Verification for semantic equivalence in mathematical reasoning tasks, and (4) LLM-as-a-Judge. Additional enhancements include Language consistency enforcement for detecting non-Korean responses and Answer extraction for identifying final answers from reasoning chains.
@@
3.3 Advanced Evaluation Pipeline
Beyond the standard Evaluator API, HRET provides a flexible PipelineRunner for advanced customization of evaluation workflows. This interface enables multi-stage evaluations where outputs from one model serve as inputs to another, comprehensive model comparisons across different benchmarks and metrics, custom workflows with specialized processing steps for Korean-specific evaluations, and ensemble evaluations combining complementary models or methodologies.
3.4 Korean-Specific Analysis Capabilities
HRET extends beyond standard metrics by integrating a Korean-specific morphological analyzer to deliver four core analyses: (1) Overall Performance with detailed accuracy, sample counts, and prediction distributions; (2) Subset Analysis highlighting variations across task types (e.g., reading comprehension vs. cultural knowledge); (3) Lexical Diversity via Type–Token Ratio (TTR) comparisons between correct and incorrect outputs; and (4) Error Pattern Detection through keyword-based inference of frequent failure modes.
Together, these tools form a diagnostic framework tailored to Korean linguistic characteristics—highlighting morphological limitations and vocabulary gaps that standard accuracy alone cannot capture, and guiding targeted improvements in model training and evaluation.
@@ 3.5 Usage Examples
HRET provides intuitive command-line and Python API interfaces. A simple CLI example for evaluating a Korean LLM on HAE-RAE Bench:
@@
This command loads the specified dataset and model, performs string-match evaluation, and outputs results in JSON format. Additional usage scenarios and advanced functionalities are provided in Appendix D.
4 System Evaluation
4.1 Reproducibility Validation
To validate HRET’s ability to address reproducibility challenges in Korean LLM evaluation, we conducted cross-institutional experiments across different hardware configurations and backend implementations. We evaluated three representative models (Qwen2.5-3B Qwen et al. (2025), Qwen3-8B Yang et al. (2025), Llama-3.2-3B-Instruct Meta AI (2024)) on Korean benchmarks using log-likelihood evaluation across four hardware setups with varying CPU architectures, GPU types, and memory configurations.
Model Max Variation Hardware Configs
Qwen2.5-3B 0.0016 (0.56%) 4 setups
Qwen3-8B 0.0028 (1.06%) 4 setups
Llama-3.2-3B 0.0015 (0.62%) 4 setups
Table 3: Hardware robustness summary across environments (detailed table in Appendix C).
Results demonstrate remarkable consistency despite significant hardware variations. Accuracy scores remained stable across different CPU architectures, GPU types, and batch sizes (1–64), with standard deviations under 0.003. Cross-institutional experiments showed identical or near-identical scores, validating HRET’s ability to eliminate evaluation discrepancies observed in Table 1.
Hardware Configuration HAE-RAE CLIcK
AMD EPYC 74F3 + H100 0.1806 0.1343
Intel Xeon + A100-80GB 0.1788 0.1298
Intel Xeon + H200 0.1802 0.1308
Standard Deviation 0.0009 0.0023
Table 4: Backend robustness using vLLM for Qwen2.5-3B.
Backend comparison between HuggingFace and vLLM shows that while absolute scores differ due to implementation differences, results remain highly consistent within each backend ($\sigma<0.003$). Compared to the 1–10 percentage point discrepancies reported between institutions, HRET achieves hardware robustness with standard deviations under 0.003 and cross-institutional consistency when using standardized procedures.
4.2 Multi-Method Evaluation Analysis
Korean language tasks exhibit diverse characteristics that benefit from different evaluation approaches. HRET’s multi-method evaluation framework enables researchers to select the most appropriate assessment strategy for each task type, leading to more accurate performance measurements.
Table 5 compares evaluation methods across different Korean benchmarks using EXAONE-3.5-2.4B Research et al. (2024) as a case study. The results demonstrate significant performance variations depending on the evaluation method employed, with notable differences between task types.
Benchmark String-Match Math-Verify LLM-as-judge(w.gpt-4o-2024-11-20)
HRM8K 2.21 14.11 17.60
CLIcK 27.31 28.67 27.76
@@ 4.2 Multi-Method Evaluation Analysis
Table 5: Comparison of evaluation methods for EXAONE-3.5-2.4B across Korean benchmarks.
Task-Dependent Sensitivity
Mathematical reasoning (HRM8K) sees dramatic score increases—from 2.21% (String-Match) to 17.60% (LLM-as-judge)—highlighting the need for semantic evaluation over exact matching.
Evaluation Robustness
Cultural and linguistic tasks (CLIcK) remain consistent (27.31–28.67%) across methods, indicating lower sensitivity to metric choice.
Methodological Implications
By matching metric to task—semantic methods for reasoning, exact-match for comprehension—HRET balances accuracy and efficiency across diverse Korean language evaluations.“‘
4.3 Test-Time Scaling Techniques
HRET incorporates advanced test-time scaling techniques to enhance evaluation quality through multiple inference strategies. These methods are particularly valuable for Korean language evaluation, where single predictions may not fully capture model capabilities due to linguistic complexity and cultural contexts.
The framework supports three primary approaches: Best-of-N Sampling, Self-Consistency with majority voting, and Beam Search for optimal decoding paths. These techniques can be combined with various evaluation methods, including LLM-as-a-judge assessments.
Table 6 demonstrates test-time scaling effectiveness on HAERAE Reasoning Challenge (HRC), evaluating Korean reasoning across science, mathematics, and quiz domains.
Model Benchmark n=1 n=3 n=5
EXAONE-3.5-2.4B HRC 3.37 5.04 9.24
Qwen2.5-3B-Instruct 11.76 12.60 13.44
Table 6: Test-time scaling results with LLM-as-a-judge(w. gpt-4o-2024-11-20) on HRC.
Results show substantial performance gains through increased sampling, with EXAONE-3.5-2.4B achieving 2.7x improvement at n=5 versus single-shot evaluation. This validates test-time scaling effectiveness for complex Korean reasoning tasks requiring nuanced understanding of context and cultural references.
4.4 Extensibility
Our registry-based design uses simple Python decorators to enable seamless extension. New Models inherit from BaseModel and add @register_model; new Datasets inherit from BaseDataset and add @register_dataset; new Evaluators inherit from BaseEvaluator and add @register_evaluator; and new Test-time scaling methods inherit from BaseScalingMethod and add @register_scaling_method. All decorated classes are auto-registered at runtime, allowing users to plug in new components with zero additional configuration.
To foster community-driven development, HRET actively welcomes external contributions and is committed to continuous maintenance and expansion. This collaborative approach ensures the framework evolves with the Korean LLM landscape while maintaining backward compatibility for existing users.